For all x>2, log(x-2)+ log x = ?
A) log (-2)
B) log(2x-2)
C) log (x^2-2x)
D) log(x-2/x)
E) log( x/x-2)

Answer by lwsshak3(11628)

For all x>2, log(x-2)+ log x = ?
Place under a single log using multiplication rule for logs
log(x-2)+ log x=log[x(x-2)]
=log(x^2-2x)
C is the correct answer.